# sample.bib

### v4

@@ 5 Ablation Study
In this section, we answer the following questions:
@@
    "How do the different components of RATE affect model performance in memory-intensive environments?"— Question 1..
@@
    "What is the benefit of using the curriculum learning? Is its use necessary?"— Question 2..
@@
    "Why do you need an MRV and what is its best configuration?"— Question 3..
Question 1.
@@
Investigating the impact of RATE components. To study the influence of memory embeddings on RATE model performance, we conducted the following experiment: for the RATE-3 and RATE-5 models trained for the T-Maze environment (with context $K=30$ on $N=3$ and $N=5$ segments, respectively), we replaced pre-trained memory embeddings with random noise vectors during inference ( Figure 5 (b)). When using random noise instead of memory embeddings, SR $=50\%$, indicating that the agent successfully reaches the junction, but only takes a turn in one direction regardless of the clue at the beginning of the corridor. Thus, we can conclude that information about is contained in memory embeddings, while the rest of the actions are shaped by the transformer parameters.
Using the Minigrid.Memory environment, the effect of each of the RATE components on the performance of the model was studied (Figure 5 (d)). So, the inference results shows that the use of memory embeddings and hidden states caching in conjunction with MRV makes the greatest contribution to the performance of the model. Using only cached hidden states or memory embeddings brings a significantly more modest contribution, albeit a positive one.
Question 2.
Curriculum learning. In the T-Maze environment, the number of actions at the junction relates to the number of actions when moving straight along the corridor as $\frac{1}{L}$ and tends to 00 as $L$ increases. In this case, due to the severe imbalance of predicted actions, RATE starts to suffer with predicting rare actions at the junction, but still shows better results compared to DT. This situation can be remedied through curriculum learning. Curriculum learning (CL) is a technique in which a model is trained on examples of increasing difficulty. During the RATE training at T-Maze, it was found that, despite the fact that RATE is able to solve the task with the standard training procedure, using CL allows to get even better results, which is reflected in Figure 5 (c).
Question 3.
@@
Memory Retention Valve architecture ablation. Without MRV in the T-Maze environment at corridor lengths of $L\gg K$, the performance of the RATE model decreased with each segment processed at inference, resulting in SR $=50$% on long trajectories. For example, in the T-Maze task, the important information to be remembered goes into memory embeddings when processing the first segment of the sequence, and then it must be retrieved when making decisions on the last segment. At the same time, due to the recurrent structure of the architecture, memory embeddings continue to be updated during the processing of intermediate segments when no new information needs to be memorized, causing important information from memory embeddings to leak out. To retain important information in memory embeddings, the Memory Retention Valve (MRV) mechanism was added to the architecture. The $5$ different schemes detailed in Appendix E were considered to implement MRV:
@@
    MRV-CA-2: cross-attention-based MRV which uses an attention mechanism to control the updating of memory embeddings and which is used in the work as the main mechanism.
@@
    MRV-CA-1: uses the same mechanism as MRV-CA-2 but the updated memory embeddings $M_{n+1}$ are fed to Query, and the incoming $M_{n}$ are fed to Key and Value.
@@
    MRV-G: gated MRV which uses a gating mechanism similar to the one used in GTrXL [parisotto2020stabilizing].
@@
    MRV-GRU: uses a GRU [gru] block to process updated $M_{n}$ with hidden states.
@@
    MRV-LSTM: uses a LSTM [hochreiter1997long] block to process updated $M_{n}$ with cached states.
@@
Inference corridor length
Model ($K_{eff}=150$) $150$ $360$ $600$ $900$
RATE w/o MRV $\ \ 1.00\pm 0.00$ $\ \ 0.66\pm 0.08$ $\ \ 0.65\pm 0.07$ $\ \ 0.61\pm 0.07$
RATE (MRV-CA-2) 1.00 ± $\pm\ 0.00$ 0.95 ± $\pm\ 0.05$ 0.90 ± $\pm\ 0.07$ 0.90 ± $\pm\ 0.07$
RATE (MRV-G) 0.86 ± $\pm\ 0.07$ 0.77 ± $\pm\ 0.08$ 0.66 ± $\pm\ 0.07$ 0.65 ± $\pm\ 0.08$
RATE (MRV-GRU) 0.99 ± $\pm\ 0.01$ 0.74 ± $\pm\ 0.07$ 0.56 ± $\pm\ 0.11$ 0.55 ± $\pm\ 0.12$
RATE (MRV-LSTM) 0.85 ± $\pm\ 0.06$ 0.64 ± $\pm\ 0.10$ 0.51 ± $\pm\ 0.11$ 0.47 ± $\pm\ 0.11$
RATE (MRV-CA-1) 0.51 ± $\pm\ 0.01$ 0.51 ± $\pm\ 0.01$ 0.49 ± $\pm\ 0.02$ 0.49 ± $\pm\ 0.01$
The best results in thein Table 3 are obtained using MRV-CA-2. This configuration is used throughout the work and is denoted simply as MRV.
@@ 6.1 Transformers for Reinforcement Learning
Transformers have found application in various areas of RL [li2023a, agarwal2023transformers]: online RL [parisotto2020stabilizing, lampinen2021towards, esslinger2022dtqn, zheng2022online, melo2022transformers, team2023human, pramanik2023recurrent], offline RL [chen2021decision, janner2021offline, reed2022gato, lee2022multi, jiang2023efficient], and model-based RL model [chen2022transdreamer, micheli2023transformers, robine2023transformerbased]. The use of transformers as a general policy for many environments is also being explored [melo2022transformers, reed2022gato, lee2022multi]. In our work, we consider the formulation of an offline model-free RL, where learning is formulated as a sequence modeling problem [chen2021decision]. Prominent representatives of such models are [chen2021decision, janner2021offline, lee2022multi, jiang2023efficient], although planning in latent space [jiang2023efficient] is considered, which can be seen as modeling of the environment. Moreover,  [janner2021offline, jiang2023efficient] are specialized for control tasks with vector observations and do not generalize to environments with observations in the form of images. Therefore, we consider the Decision Transformer [chen2021decision], which has no memory mechanism, as the main baseline for comparison.
@@ 6.2 State Space Models for Reinforcement Learning
Recently, State Space Models (SSMs) [gu2021efficiently] have shown significant success in sequence modeling, particularly in offline RL [ota2024decision, bar2023decision, cao2024mamba, gu2023mamba]. In Decision S4 (DS4) [bar2023decision], sequence modeling is executed using S4 layers within the framework of offline RL, whereas Decision Mamba (DMamba) [ota2024decision] utilizes the most recent Mamba [gu2023mamba] sequence model instead of causal self-attention. Mamba Decision Maker (MambaDM) [cao2024mamba] integrates the unique features of SSMs to effectively combine local and global features with Global-local fusion Mamba (GLoMa) module.
@@ 6.3 Memory in Transformers
There are many ways to implement the memory mechanism for transformers [dai2019transformerxl, rae2019compressive, ding2020ernie, wu2020memformer, lei2020mart, wu2022memorizing, bulatov2022rmt]. In Transformer-XL [dai2019transformerxl], it is proposed to split a long data sequence into segments and to access past segments at the expense of memory, but to ignore very distant segments. This increases the effective length of the context. The Compressive Transformer [rae2019compressive] uses compressed memory, allowing previous versions of memory to be compressed rather than discarded as in Transformer-XL. ERNIE-Doc [ding2020ernie] suggests using the retrospective feed mechanism and the enhanced recurrence mechanism. Memformer [wu2020memformer] uses external dynamic memory to encode and retrieve past information. MART [lei2020mart] extends this idea by adding a memory update mechanism similar to a recurrent neural network [hochreiter1997long, cho2014properties]. The Memorizing Transformer [wu2022memorizing] proposes to store the internal representations of past inputs. The Recurrent Memory Transformer [bulatov2022rmt] includes additional read and write memory tokens at each segment’s beginning and end. This method allows the effective context to be expanded to over 1 million tokens [zhu2020deformable]. Our work is the first attempt to adapt transform models with memory to generate agent actions in various reinforcement learning tasks in the Offline-RL setting. An Adaptive Agent (AdA) [pmlr-v202-bauer23a] uses memory architectures to store and employ information previously acquired by the agent. The default memory architecture is Transformer-XL [dai2019transformerxl] with normalization before each layer [parisotto2020stabilizing], and the use of gating on the feedforward layers [shazeer2020glu] to stabilize training. We also use Transformer-XL in our work but refrain from using additional modifications to stabilize training. Another distinctive feature of using a transformer in AdA, as opposed to DT, is that pixel observations, past actions, past rewards, and additional information are not tokenized separately but are combined into a single vector that feeds the transformer. The transformer itself predicts not only actions but also value function values.
@@ 7 Conclusion
In this paper, we propose the Recurrent Action Transformer with Memory (RATE), an autoregressive transformer model for offline reinforcement learning problems that exploits the mechanisms of recurrent memory and Memory Retention Valve (MRV) that control memory updating. In extensive experiments on environments such as ViZDoom-Two-Colors, Memory Maze, Minigrid.Memory and T-Maze, we have shown that RATE outperforms models without recurrent memory. We also show that the proposed model achieves better or comparable results against SOTA Mamba-based models in Atari and MuJoCo environments, confirming its versatility. We thoroughly investigated the influence of memory mechanisms on the model’s performance and showed that the model does indeed use them when making a decision. This method shows great potential for tackling complex tasks involving long sequences, particularly in robotics where training agents on pre-collected datasets is highly beneficial.
@@
When it comes to modeling short sequences where all the data fits into the context of a transformer, the RATE model is of limited use. In these cases, the effectiveness of the memory mechanisms is limited because there is no need for processing long dependencies, which allows the use of a common transformer.
Impact Statements.
In this paper, we present research intended to further the development of machine learning. There are many potential societal effects of our work, but we believe that none need special emphasis in this discussion.
[…]
Appendix A Environments
A.1 Memory-intensive environments
@@
We used a modified ViZDoom-Two-Colors environment from [memup] to assess the model’s memory abilities. The agent initially having $100$ hit points (HP) is placed in a room without inner walls filled with acid. At each step in the environment, the agent loses a fixed amount of health ($5/32$ HP per step). In the center of the environment, there is a pillar of either green or red color, which disappears after $45$ environment steps. Throughout the environment, objects of two colors (green and red) are generated. When the agent interacts with an object of the same color as the pillar, it gains an increase in health of $+25$ and a reward of $+1$. When the agent interacts with an object of the opposite color, it loses a similar amount of health. The agent receives an additional reward of $+0.01$ for each step it survives. The episode ends when the agent has zero health. Thus, the agent needs to remember the color of the pillar to select items of the correct color, even if the pillar is out of sight or has disappeared. The agent does not receive information about its current health or rewards, as these observations essentially convey the same information as the color of the pillar but persist beyond step $45$. We collected a dataset of $5000$ trajectories of 90 steps in length using a trained A2C agent [DBLP:journals/corr/abs-1904-01806] (an agent trained with a non-disappearing pillar). When collecting trajectories, to ensure that the agent saw the pillar before it disappeared, the agent always appeared facing the pillar in the same place – midway between the pillar and the nearest wall. In order to successfully complete this task, the agent needs to remember the color of the pillar. This environment tests the long-term memory mechanism, since the agent needs to retain information about the pillar for a time much longer than the pillar has been in the environment. Using only short-term memory and, for example, collecting the next item of the same color as the previous collected item, it will not be possible for the agent to survive for a long time, as this policy is extremely unstable. This is due to the fact that in the training dataset the agent occasionally makes a mistake and picks up an object of the opposite color. Thus, irrelevant information about the desired color may enter the transformer context and the agent will start collecting items of an opposite color, which will quickly lead to a failure.
@@
To investigate agent’s long-term memory on very long environments (the inference trajectory length is much longer than the context length $K$) we used a modified version of the T-Maze environment [ni2023when]. The agent’s objective in this environment is to navigate from the beginning of the T-shaped maze to the junction and choose the correct direction, based on a signal given at the beginning of the trajectory using four possible actions $a\in\{left,up,right,down\}$. This signal, represented as the $clue$ variable and equals to zero everywhere except the first observation, dictates whether the agent should turn up ($clue=1$) or down ($clue=-1$). Additionally, a constraint on the episode duration $T=L+2$, where the maximum duration is determined by the length of the corridor $L$ to the junction, adds complexity to the problem. To address this, a binary flag, represented as the $flag$ variable, which is equal to $1$ one step before the junction and 00 otherwise, indicating the arrival of the agent at the junction, is included in the observation vector. Additionally, a noise channel is added to the observation vector, with random integer values from the set $\{-1,0,+1\}$. The observation vector is thus defined as $o=[y,clue,flag,noise]$, where $y$ represents the vertical coordinate. The reward $r$ is given only at the end of the episode and depends on the correctness of the agent’s turn at the junction, being $1$ for a correct turn and 00 otherwise. This formulation deviates from the traditional Passive T-Maze environment [ni2023when] (different observations and reward functions) and presents a more intricate set of conditions for the agent to navigate and learn within the given time constraint.
Memory Maze.
In this first-person view 3D environment [pasukonis2022memmaze], the agent appears in a randomly generated maze containing several objects of different colors at random locations. The agent’s task is to find an object of the same color in the maze as the outline around its observation image. After the agent finds an object of the desired color and steps on it, the color of the outline changes and the agent must find another object. The agent receives a $+1$ reward for stepping on the correct object. Otherwise, he receives no reward. The duration of an episode is a fixed number and is equal to $1000$. Thus, the agent’s task is to find as many objects of the desired color as possible in a limited time. The agent’s effectiveness in this environment depends on its ability to memorize the structure of the maze and the location of objects in it in order to find the desired objects faster.
Minigrid.Memory.
Minigrid.Memory [MinigridMiniworld23] is a 2D grid environment designed to test an agent’s long-term memory and credit-assignment [ni2023when]. The environment map is a T-shaped maze with a small room with an object inside it at the beginning. The agent appears at a random coordinate in the maze. The agent’s task is to reach the room with the object and memorize it, then reach the junction at the end of the maze and make a turn in the direction where the same object is located as in the room at the beginning of the maze. A reward $r_{t}=1-0.9\times\frac{t}{T}$ is given for success, and 00 for failure. The episode ends after any agent turns at a junction or after a limited amount of time (95 steps) has elapsed. The agent’s observations are limited to a $3\times 3$ size frame.
A.2 Standard benchmarks
Atari games.
For the Atari game environments [bellemare2013arcade], we used the same dataset as in DT, namely the DQN replay dataset with grayscale state images [agarwal2020optimistic]. This dataset contains $500$ thousand of the $50$ million steps of an online DQN agent for each game. We use the following set of games: SeaQuest, Breakout, Pong and Qbert.
MuJoCo.
Despite the fact that memory is not explicitly used in decision making in control environments like MuJoCo [d4rl], we conducted additional experiments in this environment to compare with DT. For the continuous control tasks, we selected a standard MuJoCo locomotion environment and a set of trajectories from the D4RL benchmark [d4rl]. Since we chose DT and TAP as the main models for comparison on this data, we focused on the environments used in both works (HalfCheetah, Hopper, and Walker). We used three different dataset settings: 1) Medium1111 million timesteps generated by a “medium” policy that achieves about a third of the score of an expert policy; 2) Medium-Replay – the replay buffer of an agent trained with the performance of a medium policy (about $200$k–$400$k timesteps in our environments); 3) Medium-Expert1111 million timesteps generated by the medium policy concatenated with 1 million timesteps generated by an expert policy. The scores for the MuJoCo experiments are normalized such that 100 represents an expert policy, following the benchmark protocol outlined in [d4rl]. The performance metrics for CQL and TAP are reported from the TAP paper [jiang2023efficient], and for DT from the DT paper [chen2021decision], as they use the same dataset and evaluation protocol.
@@ A.2 Standard benchmarks
Minigrid.Memory Image Continuous Discrete
@@
Atari Image Continuous* Discrete

### v5

@@ 5 Ablation Study
In this section, we answer the following research questions (RQs) to evaluate the impact of memory on model performance:
@@
    "How do the different components of RATE affect model performance in memory-intensive environments?"— RQ 1..
@@
    "What is the upper-bound estimate of the performance of the RATE model?"— RQ 2..
@@
    "Why do you need an MRV and what is its best configuration?"— RQ 3..
Figure 6: Results of replacing RATE memory tokens and cached hidden states with white noise during inference in ViZDoom-Two-Colors.
RQ 1.
@@
Investigating the impact of RATE components. To study the influence of memory embeddings on RATE model performance, we conducted the following experiment: for the RATE model trained for T-Maze (with context $K=30$ on $N=3$ segments), we replaced pre-trained memory embeddings $M$ with random noise vectors during inference (see Figure 7).
Using random noise instead of memory embeddings, SR $=50\%$ in the T-Maze, indicating the agent reaches the junction but turns in only one direction regardless of the clue. Thus, we conclude that clue information is in memory embeddings, while the rest of the actions are shaped by transformer parameters.
During RATE inference in ViZDoom-Two-Colors environment with replacement of different components of memory RATE mechanisms (memory embeddings and cached hidden states) with noise (see Figure 6), it is found that for this environment the caching of hidden states has the largest contribution, because with its noise performance drop is the largest. Thus, in sparse environments, memory embeddings contribute the most, while in continuous environments, hidden states caching is most impactful.
Figure 7: Results of replacing RATE memory tokens with white noise curing inference in T-Maze.
RQ 2.
[…]
Table 5: Comparison of OracleDT with RATE. OracleDT determines the upper-bound estimate for the maximum reward that can be obtained using RATE in the environment. SR – Success Rate.
T-Maze
OracleDT
($K=90$)
DT (Chen et al., 2021)
($K=90$)
RATE
($K_{eff}=3\times 30=90$)
SR ($T=90$) 1.0 $\pm$ 0.0 1.0 $\pm$ 0.0 1.0 $\pm$ 0.0
SR ($T=480$) 1.0 $\pm$ 0.0 0.5 $\pm$ 0.0 0.90 $\pm$ 0.07
SR ($T=900$) 1.0 $\pm$ 0.0 0.5 $\pm$ 0.0 0.90 $\pm$ 0.07
ViZDoom-Two-Colors
Reward[Total] 56.5 $\pm$ 0.8 24.8 $\pm$ 1.4 41.5 $\pm$ 1.0
Reward[Reds] 55.3 $\pm$ 1.6 7.2 $\pm$ 0.4 38.2 $\pm$ 5.1
Reward[Greens] 57.2 $\pm$ 0.5 42.3 $\pm$ 3.3 44.7 $\pm$ 5.8
RQ 3.
@@
Memory Retention Valve architecture ablation. Without MRV in the T-Maze environment at corridor lengths of $L\gg K$, the performance of the RATE model decreased with each segment processed at inference, resulting in almost SR $=50$% on long trajectories (see Table 6). For example, in the T-Maze task, the important information to be remembered goes into memory embeddings when processing the first segment of the sequence, and then it must be retrieved when making decisions on the last segment. At the same time, due to the recurrent structure of the architecture, memory embeddings continue to be updated during the processing of intermediate segments when no new information needs to be memorized, causing important information from memory embeddings to leak out. To retain important information in memory embeddings, MRV mechanism was added to the architecture. We considered the five different schemes detailed in subsection F.3 to implement MRV:
@@
    MRV-CA-1: cross-attention-based MRV which uses an attention mechanism to control the updating of memory embeddings. The updated memory embeddings $M_{n+1}$ are fed to Query, and the incoming $M_{n}$ are fed to Key and Value.
@@
    MRV-CA-2: uses the same mechanism as MRV-CA-1 but the incoming memory embeddings $M_{n}$ are fed to Query, and the updated $M_{n+1}$ are fed to Key and Value.
@@
    MRV-G: gated MRV which uses a gating mechanism similar to the one used in GTrXL (Parisotto et al., 2020).
@@
    MRV-GRU: uses a GRU (Chung et al., 2014) block to process updated $M_{n}$ with hidden states.
@@
    MRV-LSTM: uses a LSTM (Hochreiter & Schmidhuber, 1997) block to process updated $M_{n}$ with cached states.
@@
Inference corridor length Model ($K_{eff}=30\times 5=150$) $150$ $360$ $600$ $900$ RATE w/o MRV 1.00 ± $\pm\ 0.00$ 0.66 ± $\pm\ 0.08$ 0.65 ± $\pm\ 0.07$ 0.61 ± $\pm\ 0.07$ RATE (MRV-CA-2) 1.00 ± $\pm\ 0.00$ 0.95 ± $\pm\ 0.05$ 0.90 ± $\pm\ 0.07$ 0.90 ± $\pm\ 0.07$ RATE (MRV-G) 0.86 ± $\pm\ 0.07$ 0.77 ± $\pm\ 0.08$ 0.66 ± $\pm\ 0.07$ 0.65 ± $\pm\ 0.08$ RATE (MRV-GRU) 0.99 ± $\pm\ 0.01$ 0.74 ± $\pm\ 0.07$ 0.56 ± $\pm\ 0.11$ 0.55 ± $\pm\ 0.12$ RATE (MRV-LSTM) 0.85 ± $\pm\ 0.06$ 0.64 ± $\pm\ 0.10$ 0.51 ± $\pm\ 0.11$ 0.47 ± $\pm\ 0.11$ RATE (MRV-CA-1) 0.51 ± $\pm\ 0.01$ 0.51 ± $\pm\ 0.01$ 0.49 ± $\pm\ 0.02$ 0.49 ± $\pm\ 0.01$
The best results in Table 6 were obtained using cross-attention scheme (MRV-CA-2), in which we fed $M_{n}$ memory tokens from the transformer input to the query and $M_{n+1}$ memory tokens from the transformer output to the key and value. This configuration is used throughout the work and is denoted simply as MRV. This configuration acts as an effective gating mechanism to prevent the loss of important information in prolonged sparse environments, which is reflected in significantly better results for RATE in the T-Maze environment.
@@ 6.1 Transformers for Reinforcement Learning
Transformers have found application in various areas of RL (Agarwal et al., 2023; Li et al., 2023): online RL (Parisotto et al., 2020; Lampinen et al., 2021; Esslinger et al., 2022; Melo, 2022; Zheng et al., 2022; Pramanik et al., 2023; Team et al., 2023), offline RL (Chen et al., 2021; Janner et al., 2021; Lee et al., 2022; Reed et al., 2022; Jiang et al., 2023), and model-based RL (Chen et al., 2022; Micheli et al., 2023; Robine et al., 2023). The use of transformers as a general policy for many environments is also being explored (Lee et al., 2022; Melo, 2022; Reed et al., 2022). In our work, we consider the formulation of an offline model-free RL, where learning is formulated as a sequence modeling problem (Chen et al., 2021). Prominent representatives of such models are (Chen et al., 2021; Janner et al., 2021; Lee et al., 2022; Jiang et al., 2023), although planning in latent space (Jiang et al., 2023) is considered, which can be seen as modeling of the environment. Moreover,  (Janner et al., 2021; Jiang et al., 2023) are specialized for control tasks with vector observations and do not generalize to environments with observations in the form of images. Therefore, we consider the Decision Transformer (Chen et al., 2021), which has no memory mechanism, as the main baseline for comparison.
6.2 Recurrent Neural Networks for Reinforcement Learning
For long input sequences, recurrent networks may have computational advantages over transformers. The RNNs recurrent unit maintains a hidden state, which is essentially a form of memory that is important for solving POMDPs. In Decision LSTM (DLSTM) (Siebenborn et al., 2022) in DT the transformer is replaced by an LSTM unit (Hochreiter & Schmidhuber, 1997).
@@ 6.3 State Space Models for Reinforcement Learning
Recently, State Space Models (SSMs) (Gu et al., 2021) have shown significant success in sequence modeling, particularly in offline RL (Bar-David et al., 2023; Cao et al., 2024; Gu & Dao, 2023; Ota, 2024). In Decision S4 (DS4) (Bar-David et al., 2023), sequence modeling is executed using S4 (Gu et al., 2021) layers within the framework of offline RL, whereas Decision Mamba (DMamba) (Ota, 2024) utilizes the most recent Mamba (Gu & Dao, 2023) sequence model instead of causal self-attention. Mamba Decision Maker (MambaDM) (Cao et al., 2024) integrates the unique features of SSMs to effectively combine local and global features with Global-local fusion Mamba (GLoMa) module.
@@ 6.4 Memory in Transformers
There are many ways to implement the memory mechanism for transformers (Bulatov et al., 2022; Dai et al., 2019; Ding et al., 2020; Lei et al., 2020; Rae et al., 2019; Wu et al., 2020; 2022). In Transformer-XL (TrXL) (Dai et al., 2019), it is proposed to split a long data sequence into segments and to access past segments at the expense of memory, but to ignore very distant segments. This increases the effective length of the context. The Compressive Transformer (Rae et al., 2019) uses compressed memory, allowing previous versions of memory to be compressed rather than discarded as in TrXL. ERNIE-Doc (Ding et al., 2020) suggests using the retrospective feed mechanism and the enhanced recurrence mechanism. Memformer (Wu et al., 2020) uses external dynamic memory to encode and retrieve past information. MART (Lei et al., 2020) extends this idea by adding a memory update mechanism similar to a recurrent neural network (Cho et al., 2014; Hochreiter & Schmidhuber, 1997). The Memorizing Transformer (Wu et al., 2022) proposes to store the internal representations of past inputs. The Recurrent Memory Transformer (RMT) (Bulatov et al., 2022) includes additional read and write memory tokens at each segment’s beginning and end. This method allows the effective context to be expanded to over 1 million tokens (Zhu et al., 2020).
An Adaptive Agent (AdA) (Bauer et al., 2023) uses memory architectures to store and employ information previously acquired by the agent. The default memory architecture is TrXL with normalization before each layer (Parisotto et al., 2020), and the use of gating on the feedforward layers (Shazeer, 2020) to stabilize training. We also use TrXL in our work but refrain from using additional modifications to stabilize training. Another distinctive feature of using a transformer in AdA, as opposed to DT, is that pixel observations, past actions, past rewards, and additional information are not tokenized separately but are combined into a single vector that feeds the transformer. The transformer itself predicts not only actions but also value function values.
@@ 7 Conclusion
In this paper, we propose Recurrent Action Transformer with Memory (RATE), a transformer model for offline RL that exploits memory mechanisms in the form of memory embeddings and caching of hidden states, and the Memory Retention Valve (MRV), which controls memory updating and prevents the loss of important information in sparse tasks. In extensive experiments in memory-intensive environments such as ViZDoom-Two-Colors, Memory Maze, Minigrid-Memory, and T-Maze, we have shown that RATE outperforms recurrent and transformer baselines. The proposed model interpolates and extrapolates well outside the transformer context, is able to retain important information for a long time when operating in highly sparse environments, and allows to compensate for the effect of bias in the training data.
We also show that the proposed model achieves better or comparable results to state-of-the-art Mamba-based models in Atari and MuJoCo environments, indicating that RATE is suitable for all tasks: both memory-intensive and not. We have thoroughly investigated the influence of memory mechanisms on the performance of the model and have clearly shown that the model uses them in decision making. This method shows great potential for tackling complex tasks with long sequences, especially in robotics, where training agents on pre-collected data sets is highly advantageous.
@@
Limitations of the proposed model include its inability to design $K$ and $N$ in memory-intensive environments so that all important events fall into the efficient context of $K_{eff}=K\times N$. In addition, the approach based on dividing trajectories into segments during inference does not allow for memory updates effectively at each step using a sliding window. Also, there are currently no studies of the memory capacity of the proposed model, so the practical amount of information that can be stored remains unknown.
Reproducibility Statement.
The model description is presented in section 3 (Algorithm 1 and Algorithm 2), the training procedure is presented in Appendix D, the description of the used benchmarks is presented in Appendix B, the hyperparameters are presented in Table 8, and the configurations for displaying the experimental results are presented in Table 9. The code is open-sourced and can be found in the https://sites.google.com/view/rate-model/.
Appendix A Decision Transformer
Algorithm 3 Decision Transformer
Input: $R\in\mathbb{R}^{1\times T},o\in\mathbb{R}^{d_{o}\times T},a\in\mathbb{R}^{1\times T}$
1:  $\tilde{R}\in\mathbb{R}^{T\times d}\leftarrow\texttt{Encoder}_{R}(R)$ $\ \tilde{o}\in\mathbb{R}^{T\times d}\leftarrow\texttt{Encoder}_{o}(o)$ $\ \tilde{a}\in\mathbb{R}^{T\times d}\leftarrow\texttt{Encoder}_{a}(a)$
2:  $\tau_{0..T}\leftarrow\{(\tilde{R}_{0},\tilde{o}_{0},\tilde{a}_{0}),\ldots,(\tilde{R}_{T},\tilde{o}_{T},\tilde{a}_{T})\}$
3:  $n=\texttt{random}(0,T-K)$
4:  $\hat{a}_{n}\leftarrow\texttt{Transformer}(\tau_{n..n+K})$
Output: $\hat{a}_{n}\rightarrow\mathcal{L}(a_{n},\hat{a}_{n})$
Decision Transformer (DT) (Chen et al., 2021) is an algorithm for offline RL that reduces the RL task to a sequence modeling task. In DT, the scheme of which is presented in Algorithm 3, the trajectory $\tau$ is not divided into segments as in RATE. Instead, random fragments of length $K$ are sampled from the trajectory, since originally this architecture was designed to work only with MDP. The predicted actions $\hat{a}$ are sampled autoregressively.
Appendix B Environments
B.1 Memory-intensive environments
In this section, we provide an extended description of the environments used in this paper, as well as the methodology used to collect the trajectories. Table 7 summarizes the observations type, rewards type, and actions type for each of the environments considered in this paper.
@@ B.1 Memory-intensive environments
We used a modified ViZDoom-Two-Colors environment from (Sorokin et al., 2022) to assess the model’s memory abilities. The agent initially having $100$ hit points (HP) is placed in a room without inner walls filled with acid. At each step in the environment, the agent loses a fixed amount of health ($10/32$ HP per step). In the center of the environment, there is a pillar of either green or red color, which disappears after $45$ environment steps. Throughout the environment, objects of two colors (green and red) are generated. When the agent interacts with an object of the same color as the pillar, it gains an increase in health of $+25$ and a reward of $+1$. When the agent interacts with an object of the opposite color, it loses a similar amount of health. The agent receives an additional reward of $+0.02$ for each step it survives. The episode ends when the agent has zero health. Thus, the agent needs to remember the color of the pillar to select items of the correct color, even if the pillar is out of sight or has disappeared. The agent does not receive information about its current health or rewards, as these observations essentially convey the same information as the color of the pillar but persist beyond step $45$.
We collected a dataset of $5000$ trajectories of $90$ steps in length using a trained A2C (Beeching et al., 2019) agent (an agent trained with a non-disappearing pillar). The average reward for these $90$ steps is $4.46$. When collecting trajectories, to ensure that the agent saw the pillar before it disappeared, the agent always appeared facing the pillar in the same place – midway between the pillar and the nearest wall. In order to successfully complete this task, the agent needs to remember the color of the pillar. This environment tests the long-term memory mechanism, since the agent needs to retain information about the pillar for a time much longer than the pillar has been in the environment. Using only short-term memory and, for example, collecting the next item of the same color as the previous collected item, it will not be possible for the agent to survive for a long time, as this policy is extremely unstable. This is due to the fact that in the training dataset the agent occasionally makes a mistake and picks up an object of the opposite color. Thus, irrelevant information about the desired color may enter the transformer context and the agent will start collecting items of an opposite color, which will quickly lead to a failure.
@@
To investigate agent’s long-term memory on very long environments (the inference trajectory length is much longer than the effective context length $K_{eff}$) we used a modified version of the T-Maze environment (Ni et al., 2023). The agent’s objective in this environment is to navigate from the beginning of the T-shaped maze to the junction and choose the correct direction, based on a signal given at the beginning of the trajectory using four possible actions $a\in\{left,up,right,down\}$. This signal, represented as the $clue$ variable and equals to zero everywhere except the first observation, dictates whether the agent should turn up ($clue=1$) or down ($clue=-1$). Additionally, a constraint on the episode duration $T=L+2$, where the maximum duration is determined by the length of the corridor $L$ to the junction, adds complexity to the problem. To address this, a binary flag, represented as the $flag$ variable, which is equal to $1$ one step before the junction and 00 otherwise, indicating the arrival of the agent at the junction, is included in the observation vector. Additionally, a noise channel is added to the observation vector, with random integer values from the set $\{-1,0,+1\}$. The observation vector is thus defined as $o=[y,clue,flag,noise]$, where $y$ represents the vertical coordinate. The reward $r$ is given only at the end of the episode and depends on the correctness of the agent’s turn at the junction, being $1$ for a correct turn and 00 otherwise. This formulation deviates from the traditional Passive T-Maze environment (Ni et al., 2023) (different observations and reward functions) and presents a more intricate set of conditions for the agent to navigate and learn within the given time constraint.
The dataset consists of $2000$ of trajectories for each segment of length $30$ (i.e. $6000$ trajectories for the $K_{eff}=3\times 30=90$) and consists only of successful episodes. An artificial oracle with a priori information about the environment was used to generate the dataset.
@@
Minigrid-Memory Image Sparse Discrete
@@
Action Associative Retrieval Vector
Sparse
& Discrete
Discrete Vector
Atari Image
Sparse
& Discrete
Discrete